OpenStudy (anonymous):

A is directly proportional to both b and h. A=7 when B=2 and H=7

hartnn (hartnn):

A=kBH find , k put A=7,B=2,H=7 in A=kBH

OpenStudy (anonymous):

stil there??? how do i answer this after i sub the no.???

hartnn (hartnn):

7=k*2*7 gives, k=1/2 so, A=(1/2)BH

OpenStudy (anonymous):

tnx u r really genius ahm how about this y varies inversely with both x and z. y=3 when x=4 and z=0.05

OpenStudy (anonymous):

sorry im really slow in math